The sum of #200 was deposited in the bank,which pays an interest of 7/2% per annum. Find the simple interest amount to be received after 2years.
Answer by Theo(13342)

7/2% is equal to 3.5% which is equal to a decimal equivalent of .035.

if you take 7/2% and divide it by 100, you get the same thing.

(7/2) / 100 = (7/2) * (1/100) = 7/200 = .035.

sinple interest after 2 years is equal to 2 * .035 * 200 = 14.